Given the following demand and cost function, find the prices and output in each Market and the total profit, Assume that there are only two firms in the market.

P1 = 32 – 2Q1

P1 = 22 – Q2

TC = 10 + 2Q1 + Q2


1
Expert's answer
2021-12-19T18:10:40-0500

TR1=Q1*P1=(32 – 2Q1)Q1

TR2=Q2*P2= (22 – Q2)Q2

MR1=32-4Q1

MR2= 22 – 2Q2

MC=TC=3

32-4Q1=3

Q1=7.25

22 – 2Q2=3

Q2=9.5

P1=32 – 2Q1=32-2*7.25=17.5

P2= 22 – Q2=22-9.5=12.5


TC = 10 + 2Q1 + Q2=10+2*7.25+9.5=34

TR=7.25*17.5+9.5*12.5-34=211.625